# A Modular Safety Filter for Safety-Certified Cyber-Physical Systems

### v1

@@ IV Centralized Predictive Safety Filter for Multi-Agent Mobile Robot System
where $x_{i}$ and $y_{i}$ represents the position vector $p_{i}=[x_{i},y_{i}]^{\top}$, $\theta_{i}$ is the heading, $v_{i}$ and $\omega_{i}$ are the control inputs $u_{i}=[v_{i},\omega_{i}]^{\top}$, and =$\mathcal{I}=\{1,2,...,20\}$ is an index set indicating each agent. This non-holonomic model makes the control problem more challenging than a linear double integrator.
@@
We define the safety constraints through the admissible set (Definition 1), as the minimum distance between two arbitrary agents ($d_{[i,j]}\geq\delta_{a}$) and the minimum distance between one arbitrary agent and walls ($d_{w_{[i,j]}}\geq\delta_{w}$), described by (IV). A graphical visualization of these constraints for two agents as well as the inertial frame is depicted in Fig. 3. Note that constraint set (IV) is user-defined and only specifies the admissible set. The safe set is also dependent on the velocities, and actuator limitations which are implicitly considered by the MSF.
@@
Note that the admissible set is explicitly defined via (IV), and the safe set is defined implicitly via the optimization problem (III), however, the final safe set still must be explicitly defined. For a multi-agent mobile system, we define the final safe set as a rest point for each robot when the velocity is equal to zero and the distance between two arbitrary robots is greater than a positive threshold (15d-15h). To avoid the collision, we define constraints over the backup trajectories for each agent as a minimum distance between the backup trajectories over the prediction horizon. The safety filter must be able to find safe backup trajectories that do not collide and have zero velocity at the end of the prediction horizon. We emphasize that this design is not case-dependent for mobile robots, it can be applied to similar scenarios like a group of aerial robots or any multi-agent system that has an equilibrium point. A circular reference trajectory with constant radius, $r_{0}$, and constant angular velocity, $\omega_{0}$, is defined as the formation task for the multi-agent mobile robot system as follows:
@@
where $|\mathcal{I}|$ is the cardinality of $\mathcal{I}$. The centralized safety filter for the multi-agent mobile robot system is defined as follows:
$U_{s}=\operatorname*{arg\,min}_{U_{n\in\mathcal{N}}^{k}}||U_{a}-U_{0}^{k}||^{2}_{2}$ (15a)
@@
$X_{n+1}^{k}=f_{d}(X_{n}^{k},U_{n}^{k}),$ (15c)
$U_{n}^{k}\in\mathcal{U},$ (15d)
$X_{0}^{k}=X(k),$ (15e)
${d_{n}^{k}}_{[i,j]}\leq\delta_{a},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall n\in\mathcal{N},$ (15f)
${{d_{w}}_{n}^{k}}_{[i,j]}\leq\delta_{w},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall n\in\mathcal{N},$ (15g)
${d_{N}^{k}}_{[i,j]}\leq\delta_{a},\quad{{d_{w}}_{N}^{k}}_{[i,j]}\leq\delta_{w},\quad{v}_{N}^{k}=0,$ (15h)
where ${X}$ and ${U}$ are the stacked states and control inputs for all agents defined as $X=[x_{1},x_{2},...,x_{20}]^{\top}$ and $U=[u_{1},u_{2},...,u_{20}]^{\top}$, respectively. ${d_{n}^{k}}_{[i,j]}$ and ${{d_{w}}_{n}^{k}}_{[i,j]}$ are the distance similar to (IV) at time $k$ and $n^{th}$ prediction element. Equations (15f-15h) are defined for the safe backup trajectory to avoid collisions.
@@ V Numerical Results
Two attack scenarios are implemented, following the attack definition in section II-B and safety filter setup in section IV. Note that the settings for the safety filter are identical for both scenarios as the safety filter does not depend on the type of attack. A description of the simulation setup and safety filter can be found in Table I. This simulation takes $20$ seconds, and attacks are applied in $t\in[5,10]$. We also employed an MPC controller for the tracking problem with a predictive horizon equal to 100, representing function $g(\cdot,\cdot)$ in section II-C, respectively. Note that the prediction horizon for the safety filter is $N=3$.
@@
Centralized Model-Predictive Safety Filter
@@
$w_{0}$ $0.1\,[\frac{\text{rad}}{\text{s}}]$ $|\mathcal{I}|$ 20
@@
V-A Intelligent Attack
Let the attacker read and manipulate the control and sensor measurement signals to perform an undetectable covert attack [21], for $k\in[500,1000]$. The attack vector on the measurement signals is described as follows:
$x_{a}(k+1)=f_{d}(x_{a}(k),u_{c}(k)),$ (16)
where $x_{a}$ at $k=500$, is equal to the system’s state. The evolution of equation (16) provides a state trajectory that an anomaly detector expects to see in the cyber layer, based on the control signal provided by the central controller, $u_{c}$. The attack vector, $u_{a}$, is computed via an optimization problem whose objective is to cause a collision at the origin. The results of simulation using the CasADi toolbox [22] for three snapshots at $t=0.1[s]$, $t=6.5[s]$, $t=20[s]$ representing before, during, and after the attack period are shown in Fig.(4-6). The regular system, shown on the left, has no safety shield while the safety-certified system, shown on the right, uses the filter defined via Algorithm (1). A video of this simulation is presented here [^3] Intelligent Attack: https://www.youtube.com/watch?v=cprja-LznkI.
@@ V-A Intelligent Attack
Figure 4: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=0.1[s]$. (Before the intelligent attack)
Figure 5: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=6.5[s]$. (During the intelligent attack)
Figure 6: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=20[s]$. (After the intelligent attack)
Figure 7: The effect of the proposed modular safety filter on the first agent’s input. (Intelligent attack)
Fig. 7 displays the safety-certified and attack inputs for the first agent. For the range $t\in[0,5.49]\cup[10,20]$, the first agent remains safe since there is no modification to the input. It is important to note that for $t\in[5,5.49]$, demonstrated by the magenta area, attack input is applied, however, there is no modification as the corresponding agent remains safe. For $t\in[5.49,10]$, as depicted by the yellow area, there is a significant correction by the filter to prevent unsafe situations. During $t\in[10,20]$, after the attack period, the system successfully recovers to its normal condition, and the safety filter has zero impact.
@@ V-B False Data Injection Attack
For all agents, we consider an FDI attack where $(v,\omega)=(v_{max},0)$. The objective of this attack is to steer all the agents outside of the admissible set, $\mathcal{X}$, as shown by the black square. The results are shown in Fig. 8 at $t=5.49[s]$. The safety-certified system prevents the agents from leaving the admissible set by forcing them to stop before reaching the boundaries. Since the results of non-attacked periods are similar to those of intelligent attacks, they are not illustrated in this paper, please visit the Link [^4] FDI Attack: https://www.youtube.com/watch?v=kBO05D3sZiE.
Figure 8: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=5.49[s]$. (During the FDI attack)
@@ VI Conclusions and Future Works
This paper presents a modular safety filter for Cyber-Physical Systems designed to ensure safety at all times, including in the presence of attacks. The safety filter ensures safety regardless of whether the received control command is safe or compromised by an attack, making it effective across various attacks without any assumptions on the attack model. Demonstrating the separation of safety and performance criteria, the proposed solution allows for safety during attacks alongside any high-performance controller. This highlights the versatility of safety filters in CPS applications, especially given that constrained controllers like MPC cannot optimize all types of cost functions.
The proposed safety filter is inspired by predictive safety filters developed for learning control. This paper illustrates the effectiveness of a modular approach to the safety of CPS that can handle nonlinear and high-order systems. Depending on the characteristics of the system, alternative safety filter solutions proposed for learning control can likely be used with minor adjustments to CPS. These methods include Control Barrier Functions (CBFs) and Hamilton-Jacobi (HJ) analysis where their extensions can account for uncertain, time-delay, and stochastic settings [23, 24, 25]. Each method comes with its advantages and disadvantages, but it is worth noting that the calculation of safe sets and backup trajectories is not as straightforward in other methods as it is in predictive filters, where they are calculated implicitly with the cost of solving an on-the-fly optimization.
We emphasize that the proposed centralized safety filter can be adapted for a decentralized scenario if each agent can communicate with its neighbors or predict their behavior. Developing a decentralized version of the proposed method is our next step to enhance practicality and reduce computational complexity [26]. A critical consideration when using safety filters in a non-deterministic setting is their tendency to introduce conservatism. If an accurate model of the system is unavailable, an extremely short prediction horizon and a small final set are adopted, or if there is a significant delay, safety filters may introduce unnecessary caution as any robust constrained solution. However, as the proposed modular solution is implemented as an add-on that is unaware of the cyber controller and attacks, any conservatism in the safety filter may result in the anomaly detector falsely identifying an attack. Further work is required to establish how conservatism affects anomaly detectors in the cyber layer, how the impact of conservatism on the anomaly detector can be mitigated, and whether communication between the modules may be required.

### v2

@@ IV Modular Predictive Safety Filter for a Multi-Agent Mobile Robot System
where $x_{i}$ and $y_{i}$ represents the position vector $p_{i}=[x_{i},y_{i}]^{\top}$, $\theta_{i}$ is the heading, $v_{i}$ and $\omega_{i}$ are the control inputs $u_{i}=[v_{i},\omega_{i}]^{\top}$, and =$\mathcal{I}=\{1,2,...,20\}$ is an index set indicating each agent.
@@
We define the state constraints as the minimum distance between two arbitrary agents ($d_{[i,j]}\geq\delta_{a}$) and the minimum distance between one arbitrary agent and walls ($d_{w_{[i,j]}}\geq\delta_{w}$), described by (11). Graphical visualization of these constraints for two agents, as well as the inertial frame, is depicted in Fig. 3. Note that constraint set (11) is user-defined and only specifies state constraints. The safe set is also dependent on the velocities and actuator limitations, which are implicitly considered by the MSF.
@@
The terminal safe control invariant set must still be explicitly defined. For a multi-agent mobile system, we define it as a set of rest points where each robot has zero velocity, and the distance between any two robots exceeds a positive threshold (13d-13h). To avoid the collision, we define constraints over the backup trajectories for each agent as a minimum distance between the backup trajectories over the prediction horizon. The safety filter must be able to find safe backup trajectories that do not collide and have zero velocity at the end of the prediction horizon. We emphasize that this design is not case-dependent for mobile robots; it can be applied to similar scenarios, such as a group of aerial robots or any multi-agent system that has an equilibrium point. A circular reference trajectory with constant radius, $r_{0}$, and constant angular velocity, $\omega_{0}$, is defined as the formation task for the multi-agent mobile robot system as follows:
@@
where $|\mathcal{I}|$ is the cardinality of $\mathcal{I}$. The modular safety filter for the multi-agent mobile robot system is defined as follows:
$U_{s}=\operatorname*{arg\,min}_{U_{t}^{k}}||U_{a}-U_{0}^{k}||^{2}_{2}$ (13a)
@@
$X_{t}^{k+1}=f_{d}(X_{t}^{k},U_{t}^{k}),$ (13c)
$U_{t}^{k}\in\mathcal{U},$ (13d)
$X_{t}^{0}=X(t),$ (13e)
${d_{t}^{k}}_{[i,j]}\geq\delta_{a},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall k\in\mathcal{N},$ (13f)
${{d_{w}}_{t}^{k}}_{[i,j]}\geq\delta_{w},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall k\in\mathcal{N},$ (13g)
${d_{t}^{N}}_{[i,j]}\geq\delta_{a},\quad{{d_{w}}_{t}^{N}}_{[i,j]}\geq\delta_{w},\quad{v}_{t}^{N}=0,$ (13h)
where ${X}$ and ${U}$ are the stacked states and control inputs for all agents defined as $X=[x_{1},y_{1},\theta_{i},...,x_{20},y_{20},\theta_{20}]^{\top}$ and $U=[u_{1},v_{1},...,u_{20},v_{20}]^{\top}$, respectively. Additionally, (13c) represents the discrete form of (10) for the multi-robot system with the time step $T_{s}^{s}$ as follows:
$f_{d}(X(t),U(t))=\begin{bmatrix}x_{1}(t)\\ y_{1}(t)\\ \theta_{1}(t)\\ \vdots\\ x_{20}(t)\\ y_{20}(t)\\ \theta_{20}(t)\end{bmatrix}+T_{s}^{s}\begin{bmatrix}v_{1}(t)\cos\theta_{1}(t)\\ v_{1}(t)\sin\theta_{1}(t)\\ w_{1}(t)\\ \vdots\\ v_{20}(t)\cos\theta_{20}(t)\\ v_{20}(t)\sin\theta_{20}(t)\\ w_{20}(t)\end{bmatrix}.$ (14)
Also, ${d_{t}^{k}}_{[i,j]}$ and ${{d_{w}}_{t}^{k}}_{[i,j]}$ are the distance similar to (11) at time $t$ and $k^{th}$ prediction element. Equations (13f-13h) are defined for the backup trajectory to avoid collisions.
@@ V Numerical Results
To evaluate the effectiveness of the proposed method, two attack scenarios are implemented, following the intelligent attack definition in section II-B and safety filter setup in section IV. Note that the settings for the safety filter are identical for both scenarios, as the safety filter does not depend on the type of attack. A description of the simulation setup and safety filter can be found in Table I. This simulation takes $15$ seconds, and attacks are applied in $t\in[5,10]\,\mathrm{sec}$. We also employed an MPC controller for the tracking problem with a predictive horizon equal to $100$, representing function $u_{c}(t)=g(r(t),x_{a}(t))$ in section 6. Note that the safety filter uses a shorter prediction horizon of $N=3$, which is sufficient for safety. This demonstrates the practicality of using a short-horizon safety filter for local policies, accommodating the plant’s computational limits while leveraging a large-horizon tracking controller where resources permit.
@@
Modular Predictive Safety Filter
@@
$w_{0}$ $0.4\,[\frac{\text{rad}}{\text{s}}]$ $|\mathcal{I}|$ 20
@@
The anomaly detector in the networked controller layer is defined as:
$a(t)=\begin{cases}1&\text{if }\|X_{a}(t)-X_{c}(t)\|\geq\varepsilon,\\ 0&\text{otherwise},\end{cases}$ (15)
where $X_{c}(t)=f_{d}(X_{a}(t-1),U_{c}(t-1))$ represents the expected state after applying $U_{c}$, and $\varepsilon=10^{-6}$ is the detection threshold, which can be set as small as the solver’s numerical precision. This means that if the system’s response deviates slightly from the expected state, $X_{c}(t)$, the anomaly detector is triggered. We emphasize that (15) is used solely for simulation purposes, and any other anomaly detector can be employed.
V-A First scenario: Intelligent attack
Let the attacker read and manipulate the control and sensor measurement signals to perform an undetectable covert attack [17], for $t\in[5,10]\,\mathrm{sec}$. The attack vector on the measurement signals is described as follows:
$X_{a}(t+1)=f_{d}(X_{a}(t),U_{c}(t)),$ (16)
where $X_{a}(t)$ is equal to the system’s state $X(t)$ at $t=5\,\mathrm{sec}$. The evolution of equation (16) provides a state trajectory that the anomaly detector expects to see in the networked controller based on the control signal, $U_{c}$. On the other hand, the attack control input, $U_{a}$, is computed via an optimization problem whose objective is to cause a collision at the origin. The results of simulation using the CasADi toolbox [25] for three snapshots at $t=0.1\,\mathrm{sec}$, $t=8\,\mathrm{sec}$, $t=15\,\mathrm{sec}$ representing before, during, and after the attack period are shown in Fig.(4-6). The regular system, shown on the left, has no safety mechanism, while the safety-certified system, shown on the right, uses the filter defined via Algorithm (1). A video of this simulation is presented here [^2] Intelligent Attack: https://www.youtube.com/watch?v=kBO05D3sZiE.
@@ V-A First scenario: Intelligent attack
Figure 4: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=0.1\,\mathrm{sec}$. (Before the intelligent attack)
Figure 5: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=8\,\mathrm{sec}$. (During the intelligent attack)
Figure 6: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=15\,\mathrm{sec}$. (After the intelligent attack)
The first subplot in Fig. 7, denoted as $v$, displays the safety-certified and attack inputs for the first agent. For $t\in[0,6.14]\,\cup[10,15]\,\mathrm{sec}$, the first agent remains safe since there is no modification to the input. It is important to note that for $t\in[5,6.14]\,\mathrm{sec}$, demonstrated by the magenta area in the second subplot, attack input is applied; however, there is no modification as the corresponding agent remains safe. For $t\in[6.14,10]\,\mathrm{sec}$, as depicted by the yellow area, there is a significant correction by the filter to prevent unsafe situations. During $t\in[10,15]\,\mathrm{sec}$, after the attack period, the system successfully recovers to its normal condition, and the safety filter has zero impact. It should be noted that the anomaly detector, (15), cannot detect the intelligent attack for $t\in[5,10]\,\mathrm{sec}$, and it activates only after the attack has finished, see the third subplot in Fig. 7. This is due to the nature of the intelligent attack, which exploits system dynamics and knowledge of control input signal $U_{c}$ to generate data that the anomaly detector expects to observe, i.e., $X_{c}(t)=X_{a}(t)$, resulting in an undetectable attack.
The actual position of the system (10), denoted by $P(t)=[x_{1},y_{1},\ldots,x_{20},y_{20}]^{\top}$, and the potentially attacked position received by the networked controller $P_{a}(t)$ are illustrated in Fig 8. For $t\in[0,5]\,\mathrm{sec}$, no attack occurs, and the system operates normally, i.e., $P(t)=P_{a}(t)$. During $t\in[5,10]\,\mathrm{sec}$, the attacker attempts to drive all robots to the origin, causing a collision (as shown in the upper subplot), while generating states using (16) to simulate normal conditions and evade detection. At $t=10\,\mathrm{sec}$, when the attack ends, the anomaly detector observes a sudden jump in the plant’s states and identifies the attack; see the bottom subplot in Fig. 8 and 7. Despite this, the safety filter successfully prevents the collision by stopping the robots, as evidenced by $P(t)$ in Fig. 8.
Figure 7: Effect of the proposed modular safety filter on the first agent under an intelligent attack. Here, $v$ represents the translational velocity, and $\frac{||u-u_{s}||}{2||u_{\text{max}}||}$ denotes the normalized control input vector, and $a$ is the value of the anomaly detector. (Intelligent attack)
Figure 8: Position of agents through time: $P$ represents the actual trajectory of the plant, while $P_{a}$ denotes the trajectory received by the networked controller under attack-free and intelligent attack conditions. (Intelligent attack)
@@ V-B False Data Injection Attack
For all agents, we consider an FDI attack where $u_{a}(t)=-u_{c}(t)+[v_{max},0]^{\top}$ for all agents. The objective of this attack is to steer all the agents outside of the admissible set, $\mathcal{X}$, as shown by the black square. The results are shown in Fig. 9 at $t=5.49\,\mathrm{sec}$. The safety-certified system prevents the agents from leaving the admissible set by forcing them to stop before reaching the boundaries. Since the remaining results are similar to the intelligent attack, they are not included in this paper. For further details, please visit the Link [^3] FDI Attack: https://www.youtube.com/watch?v=cprja-LznkI
Figure 9: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=5.49\,\mathrm{sec}$. (During the FDI attack)
@@ VI Conclusions and Future Works
This paper presents a modular safety filter for cyber-physical systems designed to ensure safety at all times, including in the presence of attacks. The safety filter ensures safety regardless of whether the received control command is safe or compromised by an attack, making it effective across various attacks without any assumptions on the attack model. Demonstrating the separation of safety and performance criteria, the proposed solution allows for safety during attacks alongside any high-performance controller. This highlights the versatility of safety filters in cyber-physical system applications, especially given that constrained controllers like MPC cannot optimize all types of cost functions.
The proposed safety filter is inspired by predictive safety filters developed for learning control. This paper illustrates the effectiveness of a modular approach to the safety of CPS that can handle nonlinear and high-order systems. Depending on the system’s characteristics, alternative safety filter solutions proposed for learning control can likely be used with minor adjustments to CPS. These methods include control barrier functions and Hamilton-Jacobi analysis, where their extensions can account for uncertain, time-delay, and stochastic settings [26, 27, 28]. Each method comes with its advantages and disadvantages. Still, it is worth noting that calculating safe sets and backup trajectories is not as straightforward in other methods as in predictive filters, where they are calculated implicitly with the cost of solving an on-the-fly optimization.
We emphasize that the proposed safety filter can be adapted for a distributed scenario if each agent can communicate with its neighbors or predict their behavior. Developing a distributed version of the proposed method is our next step to enhance practicality and reduce computational complexity [29]. A critical consideration when using safety filters in a non-deterministic setting is their tendency to introduce conservatism. If an accurate model of the system is unavailable, an extremely short prediction horizon and a small final set are adopted, or if there is a significant delay, safety filters may introduce unnecessary caution as any robust, constrained solution. However, since the proposed modular solution is implemented as an add-on that is unaware of the tracking controller and attacks, any conservatism in the safety filter may cause the anomaly detector to detect an attack incorrectly. Further work is required to establish how conservatism affects anomaly detectors in the networked controller, how the impact of conservatism on the anomaly detector can be mitigated, and whether communication between the modules may be required.